OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. d = 10 A = 314 sq. units 157 sq. units 78.5 sq. units

jimthompson5910 (jim_thompson5910):

d = 10 r = d/2 = 10/2 = 5 r = 5

jimthompson5910 (jim_thompson5910):

use the formula \[\Large A = \pi*r^2\] with \[\Large \pi \approx 3.14\]

OpenStudy (anonymous):

never mind

jimthompson5910 (jim_thompson5910):

what else did you get

OpenStudy (anonymous):

78, right?

jimthompson5910 (jim_thompson5910):

you mean 78.5, or something else?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

yes, A is approximately 78.5

OpenStudy (anonymous):

thnx:)